# On the estimation of the number of components in multivariate functional principal component analysis

### v2

@@ 4 Application: Canadian weather dataset
To illustrate our simulation results, we apply the same idea on a real dataset, the Canadian weather dataset, available in the R package fda (Ramsay et al., 2023). The dataset contains daily measurements of temperature (in Celsius) and precipitation (in millimeters) for $35$ Canadian weather stations, averaged over the years 1960 to 1994. This is an example of multivariate functional data with $p=2$ defined on one dimensional domain, the temperature being the first feature and the precipitation being the second feature. We aim to estimate $M$ multivariate eigencomponents of the data using different number of univariate eigencomponents and compare the results. For that, we define two scenarios, one where $M=M_{+}$ and one where $M=M_{-}$.
We first expand the data in a B-splines basis with $10$ functions. In the first scenario, for each feature, we estimate two univariate eigencomponents ($M_{1}=2$ and $M_{2}=2$). In the second scenario, for each feature, we estimate four univariate eigencomponents ($M_{1}=4$ and $M_{2}=4$). In both scenarios, we then estimate $M=4$ multivariate eigencomponents. So, for the first scenario, $M=M_{+}=4$ and for the second scenario, $M=M_{-}=4$. Based on the simulation and Figure 1, we expect the first two multivariate eigencomponents to be similar and the other two to be unequal.
@@ 2 $4$ components $15850$ $1679$ $438$ $213$
Table 1: Estimation of the first four eigenvalues of the Canadian weather dataset using two and four univariate components for the univariate expansions.
Figure 3 presents the estimation of the eigenfunctions for the Canadian weather dataset for both scenarios. As with the eigenvalues, we notice that the first two (multivariate) eigenfunctions are approximately the same, but the other two are not equal. While we do not know the true eigenfunctions, based on simulation results, the estimated multivariate eigenfunctions from the second scenario should be closer to the truth than the estimation from the first scenario. We may explain this result as we have not estimated enough information in the univariate decomposition in the first scenario to effectively estimate the third and fourth multivariate eigencomponents. We reiterate the suggestion to estimate at most $M_{-}$ multivariate eigencomponents (which will be $M_{-}=2$ for the first scenario).
@@
Figure 3: Estimation of the first four eigenfunctions of the Canadian weather dataset using using two and four univariate components for the univariate expansions.
@@ 5 Conclusion
We have conducted a simulation study and an example on a real dataset, and the obtained results highlight two important findings. Firstly, although utilizing only a few univariate components may yield a substantial number of multivariate components, their accuracy is notably limited. Secondly, relying on the percentage of variance explained as a criterion for selecting the number of univariate components may result in an underestimation of the number of multivariate components. We, therefore, advise practitioners to exercise caution when determining the number of estimated components required in their analysis. We suggest to estimate at most $M_{-}$ multivariate components. Additionally, we strongly recommend conducting simulations that closely resemble the characteristics of the actual data to select the appropriate number of components based on the percentage of variance explained criterion.

### v3

@@ 4 Application: Canadian weather dataset
To illustrate our simulation results, we apply the same idea on a real dataset, the Canadian weather dataset (Ramsay and Silverman, 2005), available in the R package fda (Ramsay et al., 2023). The dataset contains daily measurements of temperature (in Celsius) and precipitation (in millimeters) for $35$ Canadian weather stations, averaged over the years 1960 to 1994. The data are presented in Figure 2. This is an example of multivariate functional data with $p=2$ defined on one dimensional domains, the temperature being the first feature $x^{(1)}$ and the precipitation being the second feature $x^{(2)}$. We aim to estimate $M$ multivariate principal components of the data using different numbers of univariate principal components $M_{j}$ and compare the results. We define two scenarios, one where $M=M_{+}$ and one where $M=M_{-}$.
(a) Temperature (first feature)
(b) Precipitation (second feature)
Figure 2: The daily temperature and precipitation in $35$ Canadian weather stations. Each curve represents one weather station.
We first expand the data in a B-spline basis with $10$ basis functions. In the first scenario, for each feature, we estimate two univariate principal components ($M_{1}=2$ and $M_{2}=2$). In the second scenario, for each feature, we estimate four univariate principal components ($M_{1}=4$ and $M_{2}=4$). In both scenarios, we then estimate $M=4$ multivariate principal components. So, for the first scenario, $M=M_{+}=4$ and for the second scenario, $M=M_{-}=4$. Based on the simulation and Figure 1, we expect the first two multivariate principal components to be similar and the other two to be different.
@@ 2 $4$ components $15850$ $1679$ $438$ $213$
Table 2: Estimation of the first four eigenvalues of the Canadian weather dataset using two and four univariate components.
Figure 3 presents the estimation of the eigenfunctions for the Canadian weather dataset for both scenarios. As with the eigenvalues, we notice that the first two (multivariate) eigenfunctions are approximately the same, but the other two are not. The first two principal components can be interpreted similarly in both scenarios. The first component is negative for both features, indicating that weather stations with positive scores have lower temperatures and less precipitation than average. While the first component for precipitation is relatively flat, the temperature component exhibits more variation at the beginning and end of the year. The second component contrasts winter and summer: stations with positive scores have higher temperatures and more precipitation in summer, and lower temperatures with less precipitation in winter compared to the average station. The third and fourth components differ between the scenarios. For univariate expansions with two components, the third multivariate component for temperature contrasts winter and summer, while with four univariate components, it contrasts spring and autumn. The third multivariate component for precipitation varies in magnitude depending on the number of univariate components. The fourth multivariate component for temperature is roughly flat when two univariate components are used but shows more variability with four components. For precipitation, the fourth component contrasts winter and summer when using two univariate components, but becomes flatter with a negative bump in autumn when four components are considered. These results highlight that the interpretation of the multivariate components depends on the number of univariate components used for the univariate decomposition. However, it would be preferable that the interpretation remains consistent and independent of the univariate decomposition. While we do not know the true eigenfunctions, based on simulation results, the estimated multivariate eigenfunctions from the second scenario should be closer to the truth than the estimation from the first scenario. In the first scenario, the univariate decompositions did not capture enough information to accurately estimate the third and fourth multivariate principal components. We reiterate the suggestion to estimate at most $M_{-}$ multivariate principal components (which is $M_{-}=2$ for the first scenario).
@@
Figure 3: Estimation of the first four eigenfunctions of the Canadian weather dataset using two and four univariate components.
@@ 5 Conclusion
We have conducted a simulation study and an example on a real dataset, and the obtained results highlight two important findings. Firstly, although utilising only a few univariate components may yield a substantial number of multivariate components, their accuracy is notably limited. Secondly, relying on the percentage of variance explained as a criterion for selecting the number of univariate components may result in an underestimation of the number of multivariate components. We therefore advise practitioners to exercise caution when determining the number of estimated components required in their analysis. We suggest to estimate at most $M_{-}$ multivariate components if for each univariate feature, $M_{j}$ univariate components have been estimated. Additionally, we strongly recommend conducting simulations that closely resemble the characteristics of the actual data to select the appropriate number of components based on the percentage of variance explained criterion.